OpenStudy (anonymous):

I'm trying to integrate sqrt. x^2 + 4x +5 (Integrals involving quadratics) and i'm stuck with int. sec^3 theta d(theta) .. pls. someone help me

OpenStudy (anonymous):

how did this become:

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@amistre64

hartnn (hartnn):

i think we did, sec^3 theta and you want same & exact that answer or any form will do ?

OpenStudy (anonymous):

uhmm as long as we get the solution..

hartnn (hartnn):

so, where u stuck ? we did sec^3 theta right ?

OpenStudy (anonymous):

yes i did attached a file right?

hartnn (hartnn):

your original Q is sqrt. x^2 + 4x +5 , right ? so instead of trigo. substitution, can you use standard formula for sqrt. x^2+ a^2 ??

OpenStudy (anonymous):

x=a sin theta?

hartnn (hartnn):

\( \sqrt {x^2 + 4x +5}=\sqrt{(x+2)^2+1}\) after u put u = x+2, du=dx,you get a standard integral \(\int \sqrt{u^2+1}du\)

OpenStudy (anonymous):

ok..how to integrate sqrt of u^2 +1?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (amistre64):

sec^3 sec sec^2 sec (tan^2 + 1) sec tan tan + sec integratation should be pretty straight forward, except for that sec, which has the trick of (sec+tan)/(sec+tan))

OpenStudy (anonymous):

 now wats next?